Operation Summary for April 4, 2011 - April 6, 2011 |
|
Monday April 4 The day shift began with the Tevatron (TEV) off due to an RF problem, with Operations monitoring the Recycler stash (120.3E10), with the Antiproton Source (Pbar) stacking (29.4E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off for target replacement, with the MiniBooNE and Meson FTFB T-932 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store. During the day shift, RF personnel inspected the TEV RF anode power supply. Operators prepared D0 for an access; they came out of access at 11 AM. TEV personnel conducted proton studies all afternoon. During the evening shift, Operations disabled beam to T-932 at 6:03 PM. Operations put the TEV into HEP shot setup and established store 8624 at 7:16 PM with an average initial luminosity (AIL) of 301.17E30. Tuesday April 5 The midnight shift began with Operations monitoring store 8624 (145.27E30) and the stash (147.4E10), and with Pbar stacking (20.2E10). During the midnight shift, Operations established beam to T-932 at 4:48 AM. Operations terminated store 8624 at 7:07 AM. Operators prepared D0 and CDF for an access. During the day shift, CDF and D0 completed their controlled accesses. Operations put the TEV into shot setup at 11:14 AM and established store 8627 at 12:21 PM with an AIL of 346.78E30. Operations reported that Booster RF sparking kept beam from users for an hour in the afternoon. During the evening shift, Operations disabled beam to T-932 at 6:04 PM; this completed the experiment’s run. Operations reported that between 7:30 PM and 9 PM, Booster RF sparking and a MI low-level RF (LLRF) failure kept beam from users. Wednesday April 6 The midnight shift began with Operations monitoring store 8627 (94.84E30) and the stash (322.2E10), with Pbar stacking (17.0E10), and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 8627 at 7:33 AM. Operators prepared the TEV A-E and F-Sectors for access to check out problems with a sump pump and a vacuum pump. The Plans for Wednesday The plans for today are to finish the work in F-Sector and then resume stacking, stashing, and storing. The Run Coordinator will allow Linac personnel to conduct maintenance work and allow Pbar personnel to conduct water cage work. |
|
More Information Comments and Suggestions |
| last modified 04/06/2011 email Fermilab |
| Security, Privacy, Legal | |